julia and tony play and game rock paper scissor assume that at each play they make their choices with equal probability 1/3 for each of the three moves and dependently of any previous play








A.)on any single play,what is a chance of a tie







B.) What is the chance that Julia wins at the first play








Let Random variable x-rock







The probability distribution is shown below







C.)Find P(x=1)







D.)Find P(x≥1)








1
Expert's answer
2022-02-15T17:29:14-0500

Assume missing data is:

X - 0 1 2

Frequency (f) - 4 4 1

P(X)- 4/9 4/9 1/9


Solution:

(1)

(a) P(a tie) = P(both players choose the same choice)

=P({Rock,rock}, {scissors, scissors}, {paper, paper})

=3/9 = 1/3 (since total no. of cases is 9)

(b) P(Julia wins the first play)

=6/9 (since there are 6 possibilities where tie do not occur)

=2/3

(2)

(c) "P(x=1)=4\/9"

(d) "P(x\ge1)=P(x=1)+P(x=2)=4/9+1/9=5/9"